# Learning positional encodings in transformers depends on initialization

### v3

@@ Abstract
The attention mechanism is central to the transformer’s ability to capture complex dependencies between tokens of an input sequence. Key to the successful application of the attention mechanism in transformers is its choice of positional encoding (PE). The PE provides essential information that distinguishes the position and order amongst tokens in a sequence. Most prior investigations of PE effects on generalization were tailored to 1D input sequences, such as those presented in natural language, where adjacent tokens (e.g., words) are highly related. In contrast, many real world tasks involve datasets with highly non-trivial positional arrangements, such as datasets organized in multiple spatial dimensions, or datasets for which ground truth positions are not known, such as in biological data. Here we study the importance of learning accurate PE for problems which rely on a non-trivial arrangement of input tokens. Critically, we find that the choice of initialization of a learnable PE greatly influences its ability to learn accurate PEs that lead to enhanced generalization. We empirically demonstrate our findings in three experiments: 1) A 2D relational reasoning task; 2) A nonlinear stochastic network simulation; 3) A real world 3D neuroscience dataset, applying interpretability analyses to verify the learning of accurate PEs. Overall, we find that a learned PE initialized from a small-norm distribution can 1) uncover interpretable PEs that mirror ground truth positions in multiple dimensions, and 2) lead to improved downstream generalization in empirical evaluations. Importantly, choosing an ill-suited PE can be detrimental to both model interpretability and generalization. Together, our results illustrate the feasibility of learning identifiable and interpretable PEs for enhanced generalization.
@@ 1 Introduction
Transformers commonly use ordered sequences of data, like words in a sentence. The position and order of these words are crucial to their correct interpretation. In transformers, sequences of tokens (e.g., words) are processed in parallel – not sequentially. Thus, to process tokens correctly in their intended sequence, the transformer must encode a notion of position and/or ordering of tokens. This information is encoded in its positional encoding (PE) layer – a model parameter that tags each input token with a unique location. For many common forms of data, such as natural language, text, and audio, the labeling of ground truth positional information is straightforward, since tokens are ordered sequences in 1D. This led to the original design of 1D sinusoidal PEs, which were successfully applied to natural language data, and provided general spatial information about language tokens (rather than data-specific information) (Vaswani et al., 2017). More recent investigations into the role of PE in transformers has led to a proliferation of PE schemes, each specifically designed for 1D text with different properties (Su et al., 2022; Shaw et al., 2018; Vaswani et al., 2017; Raffel et al., 2020; Li et al., 2024; Kazemnejad et al., 2023; Shen et al., 2024; Golovneva et al., 2024; Press et al., 2022). However, many interesting problems require input sequences that are not in 1D (e.g., image datasets; Li et al. (2021)), or where position information is non-trivial or not known (e.g., spatially-embedded biological data). The choice of PE significantly affects the performance of transformer models, even in simple string-based tasks (Kazemnejad et al., 2023; Ruoss et al., 2023; McLeish et al., 2024). Thus, understanding how to disambiguate and learn ideal position information from data directly would likely provide improved performance while affording increased model flexibility.
@@
If learning the optimal PE for a task can enhance downstream generalization performance, what strategies can we use to achieve this? Recent work in deep learning theory suggests various parameterizations of simple neural network models, such as weight initializations, can greatly influence their learned internal representations (Woodworth et al., 2020; Chizat et al., 2020). In particular, those studies found that weight initializations in neural networks from large-norm distributions (e.g., a normal distribution with a large standard deviation) learned random, high-dimensional representations that would “memorize” input-output relations. This learning regime is commonly-referred to as the lazy learning or neural tangent kernel (NTK) regime, as it fails to learn a structured representation of the task or input. In contrast, neural networks that were initialized from a small-norm distribution (e.g., $\mathcal{N}(0,\sigma)$ for small $\sigma$) tend to learn structured representations that accurately reflected the organization of input features and were robust to noise. This is referred to as the rich or feature learning regime (Woodworth et al., 2020; Chizat et al., 2020). (We note that choosing the initialization rank can also induce rich versus learning learning; Liu et al. (2024)). Although this theoretical framework was initially developed for simple neural networks (e.g., feed-forward networks with few hidden layers), the insights drawn from it should apply to various model architectures, including transformers (Zhang et al., 2024; Kunin et al., 2024). Given the recent interest in studying the impact of PEs on generalization, we aimed to evaluate whether the norm of PE initialization would influence the ability to properly learn a structured and accurate PE that would enhance generalization and interpretability.
@@
Here we studied how the initialization of learnable PEs in transformers influence representation learning and downstream generalization. We focused on problems containing sequences with nontrivial positioning and ordering, comparing the generalization performance of models with learned PEs to other common PE schemes (such as absolute and relative PEs). We also examined the interpretability of the learned PEs. Our primary aim was to evaluate the hypothesis that learnable PEs in the feature rich learning regime would produce interpretable position information that mirrored ground truth knowledge and improve generalization performance. We tested this hypothesis in three experiments: 1) A 2D relational reasoning task called the Latin Squares Task (LST), which is analogous to simplified Sudoku; 2) Masked prediction of a nonlinear stochastic network simulation with spatially-embedded nodes; 3) a real-world neuroscience dataset, where the task is to predict masked brain activity in 3D from spatially and heterogeneously distributed brain regions. Overall, we found that when the PE is appropriately initialized with a small norm, learnable PEs can uncover ground truth PEs, which lead to improved downstream generalization in both datasets. Note that the notion of ground truth positions is task-dependent and data-dependent. In some cases, such as in biological datasets, ground truth spatial information may be difficult to know or ambiguous. In this study, we focus on tasks in which either a ground truth is unambiguous (e.g., synthetic tasks) or in which there exists a putative ground truth (e.g., biological data with known properties). These results indicate the importance of PE choice for generalization performance, and provide insights into how to optimally discover PEs for a variety of tasks in which the ground truth PE is nontrivial or not known.
1.1 Contributions
We highlight three principal conclusions of this study.
@@ 1.1 Contributions
    Using a 2D relational reasoning task and nonlinear network simulation with known ground truth position information, we demonstrate that the ability to approximate the ground truth PE depends on initializing a PE parameter with a small norm.
@@
    We demonstrate the generality of this approach to learning nontrivial PEs in a real world 3D neuroscience dataset, where only small-norm initialized PEs learn a PE representation of brain regions that reflect brain network modularity.
@@
    We demonstrate that in all experiments, learning an accurate PE enhances downstream generalization relative to alternative and commonly-used PEs.
@@ 2 Model Architecture
For all three experiments, we used a standard encoder-only transformer architecture with four layers and embedding dimension of 160 for the LST, 64 for the stochastic nonlinear network simulation, and 64 for the fMRI data (Vaswani et al., 2017). The primary transformer manipulation was the choice of PE, and includes a mix of absolute, relative, and learnable PEs (described in ensuing sections). The formulations and definitions of common PEs are detailed in the Appendix A.1. For each experiment, we trained on 15 seeds. For simplicity of analyzing attention maps, we trained models with only a single attention head (fully-connected, bidirectional attention unless specified otherwise). However, we have included results for models with multiheaded attention (2 and 4 heads) on the LST task in Fig. A15, which reduced generalization performance. The context window for the model was either 16 tokens long for the LST (given the 4 $\times$ 4 structure of the LST paradigm), 15 tokens for the NMAR model, or 360 tokens long (for the number of brain regions in the Glasser et al. (2016) brain atlas). We used the Adam optimizer with a learning rate of 0.0001. For comparable analysis, all models were trained for a fixed number of training steps (4000 epochs for the LST; 8000 puzzles per epoch; 50k training steps for human brain data). Results reported in the main text were trained without regularization, given that we were interested in understanding the role of PE initialization in isolation (no dropout, no weight decay). However, for completeness, we include results using weight decay (AdamW, with weight decay=0.1) in the Appendix for the LST task, which yielded qualitatively similar results (Fig. A8, Table A5, Table A6). A single model/seed could be trained (4000 epochs of LST) on one NVIDIA V100 GPU in under 45 minutes.
3 Experiment 1: Relational reasoning in the Latin Square Task (LST)
@@
Figure 1: A-D) The Latin Square Task (LST). The LST involves the presentation of a 2-dimensional 4-by-4 grid populated with A) four possible symbols, blank spaces and a single target probe ("?"). The agent’s aim is to solve for the target square with the rule that each shape can only appear once in every row and column. The reasoning complexity required for an LST puzzle can be manipulated by varying the number of distinct vectors that must be integrated to solve the problem. Examples of B) 1-, C) 2-, and D) 3-vector LST puzzles (left) and their solutions (right). E) Performing the LST is intuitive in 2D. However, when formatting the task sequence for neural networks, the input must be flattened into 1D. The LST is significantly more challenging when row and column information is lost. F) The pairwise distances between token positions according to rows and columns provides the “ground truth” of how tokens relate to each other in 2D space. A successful PE would preserve the pairwise distance relationships of the ground truth encoding. G) Naively using the 1D sinusoidal PE (Vaswani et al., 2017) would provide incorrect token-wise position information, since it only considers the closeness of tokens in 1D. H) In contrast, recomputing absolute positions in 2D (with sines and cosines in the embedding dimensions) would preserve the position information of a 2D grid, even after flattening the sequence into 1D.
@@
The number of relations needed to solve a given LST puzzle can be manipulated by changing the organization of the elements within the grid (Fig 1A-D). For instance, 1-vector puzzles require integration of information across a single row or column, while 2-vector problems involve integration across a single row and column. 3-vector puzzles require information integration across three rows and/or columns. Importantly, performing the LST task requires positional information of the rows and columns. Flattening the grid to 1D without preserving the position information would significantly increase the difficulty of the LST task (e.g., see Fig. 1E). For our LST experiments, we generated 8000 training puzzles, and assigned to 1-, 2-, or 3-vector conditions. We ensured that the similarity between any generated training set puzzle was distinct from the generalization (validation) set of puzzles (i.e., the Jaccard dissimilarity > 0.8 for a test puzzle to any individual training puzzle).
3.1 PE initialization influences downstream generalization in the LST
@@
Prior work in deep learning theory suggests that choice of weight initialization can influence learned representations and downstream generalization (Jacot et al., 2020; Woodworth et al., 2020; Chizat et al., 2020). In particular, the smaller the norm of the distribution from which the neural network is initialized, the more structured the learned representations will be. We first assessed whether these intuitions would generalize to learnable PEs initialized from different Normal distributions, controlling for the standard deviation (i.e., norm). For each token embedding, we initialized a learnable PE parameter from a multivariate Normal distribution, denoted $\mathcal{N}(\mathbf{0},\mathbf{\Sigma})$, where $\mathbf{\Sigma}=\sigma\mathbf{I}$, and $\mathbf{I}$ denotes the identity matrix scaled by $\sigma$ (Fig. 2A).
@@
Figure 2: A) Transformer parameterization. We parameterize a learnable PE initialized from different distributions (i.e., $\mathcal{N}(0,\sigma\mathbf{I}$), varying $\sigma$), and study the effect initialization on downstream generalization. B) The training and validation performance across 4000 epochs for examplar initializations. C) Despite the learnability of all initialized models, the choice of $\sigma$ strongly influenced downstream generalization. Boxplots reflect variability across 15 random seeds. (See also Appendix Table A2.)
@@
We initialized PEs from distributions with $\sigma\in\{0.1,0.2,0.3,...,2.0\}$. (For each PE initialization, we trained on 15 seeds.) We trained all models for 4000 epochs, and found that all models converged (Fig. 2B). Remarkably, though the choice of $\sigma$ was the only source of variation across model parameterizations, models exhibited a wide range of generalization performance (Fig. 2C). Compared to the default initialization choice of $\sigma=1$, which had a generalization performance of 0.89 (Table A2), we found that the optimal generalization performance was produced with PEs initialized with small norms (e.g., $\sigma=0.2$; Acc=0.96; Table A2). Moreover, consistent with the NTK regime, PE’s initialized from a distribution with large $\sigma$ converged, but generalized poorly (e.g., $\sigma=2.0$, Acc=0.38; see also Table A2).
@@
While our results are consistent with the hypothesis that small norm initialized models tend to learn the most generalizable representations, we did observe a slight reduction in generalization performance for PEs initialized at $\sigma=0.1$ relative to $\sigma=0.2$. In practice, we found that the poor generalization for very small values of $\sigma$ is due to the behavior of the Adam optimizer. Specifically, vanilla SGD outperformed the generalization ability of models trained with Adam for small values of $\sigma\in\{0.01,0.05,0.1\}$ (see Fig. A7). Thus, it is important for practitioners to consider optimizers when encouraging the rich/feature training regime for very small initializations.
3.2 Learnable PEs outperform commonly-used PE schemes in the LST
@@
We next sought to benchmark the optimal learnable PE ($\sigma=0.2$, learn-0.2) relative to other standard PE schemes. These schemes included absolute 1D PE (1d-fixed, using sines and cosines; Vaswani et al. (2017)), relative PE (relative; Shaw et al. (2018)), and rotary PE (rope; Su et al. (2022)). In addition, we also evaluated performance on PE schemes that have been shown to be beneficial for algorithmic and compositional generalization tasks, including no PE with a causal attention mask (c-nope) (Kazemnejad et al., 2023) and random PE without a learnable parameter (random) (Ruoss et al., 2023). (Note that the c-nope model can implicitly learn position information due to the nature of the causal attention mask.) We also included a baseline control model without any specified PE (nope), which should not be able to learn the task in a systematic way, due to the permutation invariance of bidirectional attention in the absence of PE. Finally, we included a “ground truth” PE – an absolute 2D PE based on sines and cosines (2d-fixed) – to compare how similarly the various PEs produced attention mechanisms to this ground truth model (see Appendix A.1). (Note, that the term “ground truth” applies to any rotation of the absolute 2D PE, or any PE that preserves the original 2D row and column LST information, as depicted in Fig. 1E,F).)
@@
PE Validation acc Validation SD Training acc Training SD
2d-fixed (ground truth) 0.977 0.073 1.000 0.000
learn-0.2 0.956 0.039 1.000 0.000
relative 0.920 0.042 1.000 0.000
random 0.888 0.046 1.000 0.000
rope 0.805 0.115 1.000 0.000
1d-fixed 0.781 0.185 0.999 0.000
nope 0.334 0.020 0.509 0.002
c-nope 0.314 0.042 0.559 0.102
@@
All models, except for the nope and c-nope models converged (accuracy after 4000 epochs, c-nope=0.56, nope=0.51). Poor performance was expected for the nope model due to the lack of any explicit or implicit PE information. However, it was surprising that c-nope models neither learned nor generalized, given prior research suggesting that this PE scheme had above average performance on compositional and length generalization tasks (Kazemnejad et al., 2023). Nevertheless, as anticipated, we found the ground truth 2d-fixed model to exhibit the highest generalization performance (Table 1; Appendix Fig. A10A,D). Remarkably, the next highest performing model was the learn-0.2 model ($\sigma=0.2$), followed by the relative, random, rope, and 1d-fixed PE models, respectively. While we did not find a robust statistical difference between the ground truth 2d-fixed model (97.7%) and the learn-0.2 model (95.6%) (2d-fixed vs. learn-0.2, t-test, $t(13)=0.96$, $p=0.35$), there was a significant difference between the learn-0.2 model with the next highest-performing model (relative) ($t(13)=2.51$, $p=0.03$). Further, when applying L2 regularization (via weight decay), the generalization performance of learnable PEs was virtually indistinguishable from the ground truth 2d-fixed PE (both models generalized with 99% accuracy; Table A6; Fig. A8). Finally when performing a perturbation analysis, where we systematically injected noise to the token embeddings and evaluated downstream performance, we found that models that were most robust to noise were those initialized from a low-norm distribution (Fig. A11). Overall, these findings corroborate theoretical findings in deep learning, suggesting their applicability to learning rich transformer PEs in structured reasoning tasks.
3.3 Learnable PE models discover ground truth attention maps and positions

### v4

# Learning interpretable positional encodings in transformers depends on initialization

@@ Abstract
In transformers, the positional encoding (PE) provides essential information that distinguishes the position and order amongst tokens in a sequence. Most prior investigations of PE effects on generalization were tailored to 1D input sequences, such as those presented in natural language, where adjacent tokens (e.g., words) are highly related. In contrast, many real world tasks involve datasets with highly non-trivial positional arrangements, such as datasets organized in multiple spatial dimensions, or datasets for which ground truth positions are not known. Here we find that the choice of initialization of a learnable PE greatly influences its ability to learn interpretable PEs that lead to enhanced generalization. We empirically demonstrate our findings in three experiments: 1) A 2D relational reasoning task; 2) A nonlinear stochastic network simulation; 3) A real world 3D neuroscience dataset, applying interpretability analyses to verify the learning of accurate PEs. Overall, we find that a learned PE initialized from a small-norm distribution can 1) uncover interpretable PEs that mirror ground truth positions in multiple dimensions, and 2) lead to improved generalization. These results illustrate the feasibility of learning identifiable and interpretable PEs for enhanced generalization.
Machine Learning, ICML
@@ 1 Introduction
Transformers commonly use ordered sequences of data, like words in a sentence. The position and order of these words are crucial for their correct interpretation. In transformers, sequences of tokens (e.g., words) are processed in parallel rather than sequentially. Thus, to process tokens correctly in their intended sequence, the transformer must encode a notion of position and/or ordering of tokens. This information is encoded in its positional encoding (PE) layer – a model parameter that tags each input token with a unique location. For many common forms of data, such as natural language, text, and audio, the labeling of ground truth positional information is straightforward, since tokens are ordered sequences in 1D. This led to the original design of 1D sinusoidal PEs, which were successfully applied to natural language data, and provided general order information about language tokens (Vaswani et al., 2017). More recent investigations into the role of PE in transformers has led to a proliferation of PE schemes, each specifically designed for 1D text with different properties (Su et al., 2022; Shaw et al., 2018; Vaswani et al., 2017; Raffel et al., 2020; Li et al., 2024; Kazemnejad et al., 2023; Shen et al., 2024; Golovneva et al., 2024; Press et al., 2022). However, many interesting problems require input sequences that are not in 1D (e.g., image datasets; Li et al. (2021)), or where position information is non-trivial or not known (e.g., spatially-embedded biological data). Thus, understanding how to learn position information from data directly would likely provide improved performance while affording increased model flexibility.
@@
If learning the optimal PE for a task can enhance generalization performance, what strategies can we use to learn them? Recent work in deep learning theory suggests that model initializations can greatly influence learning dynamics (Woodworth et al., 2020; Chizat et al., 2020). In particular, those studies found that weight initializations in neural networks from large-norm distributions (e.g., a normal distribution with a large standard deviation) learned random, high-dimensional representations that would “memorize” solutions. This learning regime is commonly-referred to as the lazy learning or neural tangent kernel (NTK) regime. In contrast, neural networks that were initialized from a small-norm distribution (e.g., $\mathcal{N}(0,\sigma)$ for small $\sigma$) tend to learn structured representations that accurately reflected the organization of input features and were robust to noise. This is referred to as the rich or feature learning regime (Woodworth et al., 2020; Chizat et al., 2020). Although this theoretical framework was initially developed for simple neural networks (e.g., feed-forward networks with few hidden layers), the insights drawn from it should apply to various model architectures, including transformers (Zhang et al., 2024; Kunin et al., 2024). Given recent interest in studying the impact of PEs on generalization, we investigated whether PE initialization would influence the ability to properly learn interpretable PEs that would enhance generalization in transformers.
@@
We focused on problems containing sequences with nontrivial positioning and ordering, comparing the generalization performance of models with learned PEs to other common PE schemes (such as absolute and relative PEs). We also examined the interpretability of the learned PEs. Our primary aim was to evaluate the hypothesis that learned PEs in the feature rich learning regime would produce interpretable position information that improved generalization. We tested this hypothesis in three experiments: 1) A 2D relational reasoning task called the Latin Squares Task (LST), which is analogous to simplified Sudoku; 2) Masked prediction of a nonlinear stochastic network simulation with spatially-embedded nodes; 3) a real-world neuroscience dataset, where the task is to predict masked brain activity in 3D from spatially distributed brain regions. Overall, when the PE is appropriately initialized with a small norm, learned PEs can uncover ground truth positions that lead to improved generalization.
Summary of contributions.
@@
    Using a 2D relational reasoning task and nonlinear network simulation with known ground truth positional arrangements, we demonstrate that the ability to learn the ground truth PE depends on initializing a PE parameter with a small norm.
@@
    We demonstrate the generality of this approach to learning nontrivial PEs in a real world 3D neuroscience dataset, where only small-norm initialized PEs learn a representation of brain regions that reflect brain network clustering.
@@
    In all experiments, learning an accurate PE enhances downstream generalization relative to standard PEs.
@@ 2 Model Architecture
For all three experiments, we used a standard encoder-only transformer architecture with four layers and embedding dimension of 160 for the LST, 64 for the stochastic nonlinear network simulation, and 64 for the fMRI data (Vaswani et al., 2017). The primary transformer manipulation was the choice of PE, and includes a mix of absolute, relative, and learnable PEs (described in ensuing sections). The formulations and definitions of common PEs are detailed in the Appendix A.1. For each experiment, we trained 15 seeds. For simplicity of analyzing attention maps, we trained models with only a single attention head (fully-connected, bidirectional attention unless specified otherwise). However, we have included results for models with multiheaded attention (2 and 4 heads) on the LST task in Fig. A14. The context window for the model was either 16 tokens long for the LST (given the 4 $\times$ 4 structure of the LST paradigm), 15 tokens for the NMAR model, or 360 tokens long (for the number of brain regions in the Glasser et al. (2016) brain atlas). We used the Adam optimizer with a learning rate of 0.0001, and standard SGD in Appendix (Fig. A6). For comparable analysis, all models were trained for a fixed number of training steps (4000 epochs for the LST; 8000 puzzles per epoch; 50k training steps for human brain data). Results reported in the main text were trained without regularization, given that we were interested in understanding the role of PE initialization in isolation (no dropout, no weight decay). However, for completeness, we include results using weight decay (AdamW, with weight decay=0.1) in the Appendix for the LST task, which yielded qualitatively similar results (Fig. A7, Table A5, Table A6). A single model/seed could be trained (4000 epochs of LST) on one NVIDIA V100 GPU in under 45 minutes.
3 Experiment 1: Latin Square Task (LST)
@@
Figure 1: A-D) The Latin Square Task (LST). The LST involves the presentation of a 2D 4-by-4 grid populated with A) four possible symbols, blank spaces and a single target probe (”?”). The agent’s aim is to solve for the target square with the rule that each shape can only appear once in every row and column. The reasoning complexity required for an LST puzzle can be manipulated by varying the number of distinct vectors that must be integrated to solve the problem. Examples of B) 1-, C) 2-, and D) 3-vector LST puzzles (left) and their solutions (right). E) Performing the LST is intuitive in 2D. However, when formatting the task sequence for neural networks, the input must be flattened into 1D. The LST is significantly more challenging when row and column information is lost. F) The pairwise distances between token positions according to rows and columns provides the “ground truth” of how tokens relate to each other in 2D space. A successful PE would preserve the isometry (i.e., pairwise distance relationships) of the (x,y) grid coordinates. G) Naively using the 1D sinusoidal PE (Vaswani et al., 2017) would provide incorrect token-wise position information, since it only considers the closeness of tokens in 1D. H) In contrast, recomputing absolute positions in 2D (with sines and cosines in the embedding dimensions) would preserve the position information of a 2D grid, even after flattening the sequence into 1D.
@@
Performing the LST task requires positional information of the rows and columns. Flattening the 2D grid to 1D without preserving the position information would significantly increase the difficulty of the LST task (e.g., see Fig. 1E). (This would be like playing Sudoku in 1D.) For our LST experiments, we generated 8000 training puzzles.We ensured that the similarity between any generated training set puzzle was distinct from the generalization (test) set of puzzles (i.e., the Jaccard dissimilarity $>$ 0.8 for a test puzzle to any individual training puzzle). Simple neural network architectures, such as vanilla MLPs and bidirectional LSTMs, could not generalize well across this split (Fig. A15).
3.1 PE initialization influences LST generalization
@@
Prior work in deep learning theory suggests that the smaller the norm of the distribution from which the neural network is initialized, the more structured the learned representations will be (Jacot et al., 2020; Woodworth et al., 2020; Chizat et al., 2020). We first assessed whether these intuitions would generalize to learnable PEs initialized from different Normal distributions, controlling for the standard deviation (i.e., norm). For each token embedding, we initialized a learnable PE parameter from a multivariate Normal distribution, denoted $\mathcal{N}(\mathbf{0},\mathbf{\Sigma})$, where $\mathbf{\Sigma}=\sigma\mathbf{I}$, and $\mathbf{I}$ denotes the identity matrix scaled by $\sigma$ (Fig. 2A).
@@
Figure 2: A) Transformer parameterization. We parameterize a learnable PE initialized from different distributions (i.e., $\mathcal{N}(0,\sigma\mathbf{I}$), varying $\sigma$), and study the effect of initialization on downstream generalization. B) The training and validation performance across 4000 epochs for example initializations. C) Despite the learnability of all initialized models, the choice of $\sigma$ strongly influenced downstream generalization. Boxplots reflect variability across 15 random seeds. (See also Appendix Table A2.)
@@
We initialized PEs from distributions with $\sigma\in\{0.1,0.2,0.3,...,2.0\}$. We trained all models for 4000 epochs, and found that all models converged (Fig. 2B). Remarkably, though the choice of $\sigma$ was the only source of variation across model parameterizations, models exhibited a wide range of generalization performance (Fig. 2C). We found that the optimal generalization performance was produced with PEs initialized with small norms (e.g., $\sigma=0.2$; Acc=0.96; Table A2). Moreover, consistent with the NTK regime, PE’s initialized from a distribution with large $\sigma$ converged, but generalized poorly (e.g., $\sigma=2.0$, Acc=0.38; Fig. 2C, but see also Table A2).
@@
While our results are consistent with the hypothesis that small norm initialized models tend to learn the most generalizable representations, we did observe a slight reduction in generalization performance for PEs initialized at $\sigma=0.1$ relative to $\sigma=0.2$. In practice, we found that the poor generalization for very small values of $\sigma$ is due to the behavior of the Adam optimizer. Specifically, vanilla SGD outperformed the generalization ability of models trained with Adam for small values of $\sigma\in\{0.01,0.05,0.1\}$ (see Fig. A6). This is likely due to Adam’s adaptive learning rates, which is scaled by the magnitude of the gradient flow in Adam. Thus, it is important for practitioners to consider optimizers when encouraging the rich/feature training regime for very small initializations.
3.2 Learnable PEs outperform standard PEs in the LST
@@
We next sought to benchmark the optimal learnable PE ($\sigma=0.2$, learn-0.2) relative to other standard PE schemes. These schemes included absolute 1D PE (1d-fixed, using sines and cosines; Vaswani et al. (2017)), 1D relative PE (1d-relative; Shaw et al. (2018)), and 1D rotary PE (1d-rope; Su et al. (2022)). In addition, we also evaluated performance on PE schemes that have been shown to be beneficial for algorithmic and compositional generalization tasks, including no PE with a causal attention mask (c-nope) (Kazemnejad et al., 2023) and random PE without a learnable parameter (random) (Ruoss et al., 2023). We also included a baseline control model without any specified PE (nope), which should not be able to learn the task in a systematic way, due to the permutation invariance of bidirectional attention in the absence of PE. Finally, we included a “ground truth” PE – an absolute 2D PE based on sines and cosines (2d-fixed) – to compare how similarly various PEs produced attention maps to this ground truth model (see Appendix A.1). (Note, that the term “ground truth” applies to any rotation of the absolute 2D PE, or any PE that preserves the pairwise isometry between the row and column information in the 2D LST grid. This could include alternative absolute 2D PEs not based on sinusoids, in addition to relative PEs that preserve the isometry of the 2D grid information in the LST, as depicted in Fig. 1E,F).)
@@
PE Test acc (sd) Train acc (sd)
2d-fixed (ground truth) 0.977 (0.073) 1.000 (0.000)
learn-0.2 0.956 0.039) 1.000 (0.000)
1d-relative 0.920 (0.042) 1.000 (0.000)
random 0.888 (0.046) 1.000 (0.000)
1d-rope 0.805 (0.115) 1.000 (0.000)
1d-fixed 0.781 (0.185) 0.999 (0.000)
nope 0.334 (0.020) 0.509 (0.002)
c-nope 0.314 (0.042) 0.559 (0.102)
@@
All models, except for the nope and c-nope models converged (accuracy after 4000 epochs, c-nope=0.56, nope=0.51). Poor performance was expected for the nope model due to the lack of any explicit or implicit PE information. As anticipated, we found the ground truth 2d-fixed model to exhibit the highest generalization performance (Table 1; Appendix Fig. A9A,D). Remarkably, the next highest performing model was the learn-0.2 model ($\sigma=0.2$), followed by the 1d-relative, random, 1d-rope, and 1d-fixed PE models, respectively. While we did not find a robust statistical difference between the ground truth 2d-fixed model (97.7%) and the learn-0.2 model (95.6%) (2d-fixed vs. learn-0.2, t-test, $t(13)=0.96$, $p=0.35$), there was a significant difference between the learn-0.2 model with the next highest-performing model (1d-relative) ($t(13)=2.51$, $p=0.03$). In addition, we assessed the interaction of PE initialization with weight decay (AdamW), a commonly-used L2 regularization scheme (Loshchilov and Hutter, 2019). We empirically found that the combination of low initialization and weight decay improved generalization; performance was virtually indistinguishable from the ground truth 2d-fixed PE (both models generalized with 99% accuracy; Table A6; Fig. A7). Overall, these findings corroborate theoretical findings in deep learning, suggesting their applicability to learning interpretable PEs in transformers.
3.3 Learnable PE models learn interpretable attention maps and PEs